OpenStudy (anonymous):

find the distance between points P(2,2) and Q(7,4) to the nearest tenth. 7 10.8 5.4 4.5 thanks

OpenStudy (anonymous):

find the distance between points P(2,2) and Q(7,4) to the nearest tenth

OpenStudy (anonymous):

d = sqrt[(4-2)^2 + (7-2)^2] = sqrt[4+25] = sqrt(29) = 5.4